If 6% of a population is homozygous recessive for a mutation, what is the frequency of this mutated allele (1 mark)? Show your working out.
Answer #1

According to Hardy-Weinberg equilibrium, homozygous recessive frequency is given by q2 which is given as 6% or 0.06

Here, q is the allelic frequency for the recessive allele.

therefore, q = 0.06(1/2)

or q = 0.2449

or q = 0.25 (answer)

[Additionally, frequency of normal (wild-type, dominant) allele can also be calculated as p = 1 - 0.25 = 0.75]